If we add 2 to Connie's age it equals 1/2 the age of her brother Jack who is 12. How many years before Connie's age is exactly 1/2 of her brother Jack's age?
I know that C + 2 = 1/2(J), then C + 2 - 2 = 1/2(12) - 2, so C = 4 but I am having trouble with the rest of the equation,
I have come up with C + y = 1/2(J + y), but this does't allow me to solve for y.

Answer by nerdybill(7384)

If we add 2 to Connie's age it equals 1/2 the age of her brother Jack who is 12. How many years before Connie's age is exactly 1/2 of her brother Jack's age?
.
Let x = Connie's age
.
From:
If we add 2 to Connie's age it equals 1/2 the age of her brother Jack who is 12.
x+2 = (1/2)(12)
x+2 = 6
x = 4 (Connie's present age)
.
Let y = years before Connie's age is 1/2 her brother's
4+y = (1/2)(12+y)
8+2y = 12+y
8+y = 12
y = 4 (in four years)